40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 25 23:15:11 2013, Masayuki, Summary, Green Locking, FPMI noise caused by ARM locking fpmibd.pdf
    Reply  Wed Sep 25 23:59:29 2013, Masayuki, Summary, Green Locking, FPMI noise caused by ARM locking XARM_OLTF.pdfYARM_OLTF.pdfXARM_ITMXresponse.pngYARM_ITMYresponse.pngfree_running.pdf
       Reply  Thu Sep 26 01:49:28 2013, Masayuki, Summary, Green Locking, FPMI noise caused by ARM locking MICH_OLTF.pdfBS-RS55Q.pngETMX-RS55.pngETMY-RS55.pngplot.pdf
       Reply  Thu Sep 26 23:02:40 2013, rana, Update, LSC, FPMI noise caused by ARM locking Xarm_sweep_130926.pdflsc.pdferr.png
          Reply  Fri Sep 27 00:48:53 2013, Masayuki, Update, LSC, FPMI noise caused by ARM locking phase_badget.pngFPMI_noises.png
Message ID: 9167     Entry time: Thu Sep 26 23:02:40 2013     In reply to: 9162     Reply to this: 9168
Author: rana 
Type: Update 
Category: LSC 
Subject: FPMI noise caused by ARM locking 

Hidden in Nakano-kun's previous entries was that the phase margin of the X-Arm was only 9 degrees!! This extremely close to instability and makes for huge gain peaking. The feedback loop is increasing noise above 100 Hz rather than suppress. After some tweaks of the LSC filters we got a much more stable loop/.

So we today started to examine the sources of phase lag in the arm cavity sweeps. There were a few unfortunate choices in the XARM LSC filter bank which we tuned to get less delay.

Then I wrote a bunch of detail about how that worked, but the ELOG ate my entry because it couldn't handle converting my error signal noise plot into a thumbnail. Then it crashed and I restarted it. We also have now propagated the changes to the Y arm by copy/paste the filters and the result there is pretty much the same: low phase margin is now 38 deg phase margin. Noise is less bad.

 

Attachment 1: Xarm_sweep_130926.pdf  18 kB  | Hide | Hide all | Show all
Xarm_sweep_130926.pdf
Attachment 2: lsc.pdf  45 kB  Uploaded Fri Sep 27 00:58:48 2013  | Hide | Hide all | Show all
lsc.pdf
Attachment 3: err.png  127 kB  Uploaded Fri Sep 27 01:26:38 2013  | Show | Hide all | Show all
ELOG V3.1.3-